### v2

# Long-Context Autoregressive Video Modeling with Next-Frame Prediction

Abstract
Long-context autoregressive modeling has significantly advanced language generation, but video generation still struggles to fully utilize extended temporal contexts. To investigate long-context video modeling, we introduce Frame AutoRegressive (FAR), a strong baseline for video autoregressive modeling. Just as language models learn causal dependencies between tokens (i.e., Token AR), FAR models temporal causal dependencies between continuous frames, achieving better convergence than Token AR and video diffusion transformers. Building on FAR, we observe that long-context video modeling faces challenges due to visual redundancy. Training on long videos is computationally expensive, as vision tokens grow much faster than language tokens. To tackle this issue, we propose balancing locality and long-range dependency through long short-term context modeling. A high-resolution short-term context window ensures fine-grained temporal consistency, while an unlimited long-term context window encodes long-range information using fewer tokens. With this approach, we can train on long video sequences with a manageable token context length, thereby significantly reducing training time and memory usage. Furthermore, we propose a multi-level KV cache designed to support the long short-term context modeling, which accelerating inference on long video sequences. We demonstrate that FAR achieves state-of-the-art performance in both short- and long-video generation, providing a simple yet effective baseline for video autoregressive modeling. The code is released at https://github.com/showlab/FAR.
Index Terms:
Video Generation, Autoregressive Video Modeling, Diffusion Model.
1 Introduction
Advanced long-context autoregressive language models have demonstrated remarkable capabilities, enabling various applications that require test-time scaling, such as extended conversations [1], chain-of-thought reasoning [2, 3, 4], in-context learning [5], and retrieval-augmented generation [6]. However, video modeling has not achieved comparable progress. Recent video autoregressive modeling [7] directly adapts the paradigm of language models [8], where frames are factorized into discrete [9, 10] codes for next-token prediction (denoted as Token-AR). However, Token-AR still fails to achieve comparable quality to video diffusion transformers [11] due to the unidirectional modeling of visual tokens [12] and the irreparable information loss caused by vector quantization. On the other hand, video diffusion models [13, 14, 15, 16], which generate long videos using a progressive sliding window, struggle to effectively use earlier context.
Figure 1: Evaluation on Long Video Prediction. FAR effective exploits long video contexts and achieves accurate prediction.
In this paper, we introduce Frame AutoRegressive (FAR) model, specifically designed for video autoregressive modeling. FAR is trained using a frame-wise flow matching objective with autoregressive contexts. Unlike Token-AR, which learn causal dependencies between discrete tokens, FAR captures causal dependencies between continuous frames while still allowing full attention modeling within each frame. However, as a hybrid AR-Diffusion model [17, 18, 19], FAR also encounters a common issue observed in such models, namely, the discrepancy in observed contexts between training and inference. During training, later frames are exposed only to noised context frames due to diffusion objective, whereas inference relies on clean context frames. Recent methods [20, 21] mitigate this issue by appending a clean copy of the noised sequence during training, but this approach doubles the training cost.
To address the discrepancy of observed context, we propose training FAR with stochastic clean context. During training, we randomly replace a portion of noisy frames with clean frames and assign them a unique timestep embedding beyond the diffusion schedule to indicate representation extraction from clean context. During inference, this special embedding guides the model to effectively utilize clean context frames. We demonstrate FAR with stochastic clean context achieve same training efficiency to video diffusion transformers while achieving better convergence, served as a strong autoregressive video generation baseline.
Building on FAR, we investigate how to train on long video sequences to enable long-context video modeling. Unlike language modeling, long-context video modeling suffers from visual redundancy, as vision tokens grow significantly faster than language tokens when context increases. To address this challenge, we propose to balance the locality and long-range dependency with long short-term context modeling. Specifically, we maintain a high-resolution short-term context window to ensure fine-grained temporal consistency, while using a unlimited long-term context window with aggressive patchification to reduce redundant context tokens. This strategy enables efficient training on long video sequences with a manageable token context length. To speed up the proposed long short-term context modeling during inference, we introduce a multi-level KV cache to enhance inference efficiency.
Our contributions are summarized as follows:
  1. 1.
    We introduce FAR, an strong autoregressive video generation baseline, combined with stochastic clean context to bridge the training-inference gap in observed context.
  2. 2.
    Building on FAR, we introduce long short-term context modeling for efficient long-video training, as well as multi-level KV cache for faster inference.
  3. 3.
    FAR enables significantly lower training costs on long videos and achieves state-of-the-art performance in both short- and long-video generation.
2 Related Work
2.1 Video Generation
Video Diffusion Models. Recent advances in video generation have led to the scaling of video diffusion transformers [22, 11, 23] for text-to-video generation, resulting in superior visual quality. Pretrained text-to-video models are subsequently fine-tuned to incorporate images as conditions for image-to-video generation [24, 25, 22]. The trained image-to-video models can be utilized for autoregressive long-video generation using a sliding window [15, 16], but their ability to leverage visual context is limited by the sliding window’s size. In this work, we show that FAR achieves better convergence than video diffusion transformers for short-video generation while naturally supporting variable-length visual context.
Token Autoregressive Models. Video generation based on token autoregressive models (i.e., Token AR) aims to follow the successful paradigm of large language models. These models typically quantize continuous frames into discrete tokens [10, 26] and learn the causal dependencies between tokens using language models [7, 27]. While they achieve plausible performance, their generation quality remains inferior to that of video diffusion transformers due to information loss from vector quantization. Additionally, unidirectional visual token modeling may be suboptimal [12]. Subsequent studies have explored continuous tokens [28] without vector quantization but have not demonstrated their effectiveness in video generation. In this work, we show that FAR can learn causal dependencies from continuous frames and achieve better performance than Token AR in both short- and long-video modeling.
Hybrid AR-Diffusion Models. To leverage the strengths of both continuous latent spaces and autoregressive modeling, recent studies [19, 29, 30] have explored hybrid AR-Diffusion models. These models typically employ a diffusion objective for image-level modeling with autoregressive contexts. Hybrid AR-Diffusion models are widely applicable to both visual [19, 18, 17] and language generation [31, 32]. Recent research has also applied it in frame-level autoregressive modeling [18, 17] for video generation. However, they suffer from a training-inference discrepancy in the observed context. Some studies [20, 21] have attempted to mitigate this issue by maintaining a clean copy of the noised sequence during training, but this approach doubles the training cost. Among these methods, FAR efficiently addresses the training-inference gap through the proposed stochastic clean context, demonstrating its superior performance in long-context video modeling.
Figure 2: Illustration of FAR’s Training and Inference Pipeline. In short-video training, a portion of frames is randomly replaced with clean context frames, marked with a unique timestep embedding (e.g., -1) beyond the flow-matching scheduler. In long-video training, we adopt long short-term context modeling. A long-term context window with aggressive patchification is adopted to reduce redundant vision tokens, while a short-term context window is used to model fine-grained temporal consistency.
2.2 Long-Context Language Modeling
A straightforward approach to improving long-context ability in language modeling is to directly fine-tune the model on longer sequences. Recent work [33, 34] has explored efficient long-sequence fine-tuning with position interpolation. However, training on long videos leads to prohibitive computational costs, as vision tokens scale much faster than language tokens with increasing context length. To address this issue, we introduce long short-term context modeling to reduce visual redundancy in long-video training.
2.3 Long-Context Video Modeling
Recent advancements in video generation models have enabled their use as interactive world simulators [35, 36, 37], which require the ability to exploit long-range context and memorize the observed environment. However, existing video diffusion transformers lack effective mechanism to utilize long-range context. Although early work [38] has explored long-context video prediction, it has been limited in visual quality and long-range consistency. In this work, we introduce FAR, a efficient framework for both short- and long-context autoregressive video modeling.
TABLE I: Model Variants of FAR. We follow the model size configurations of DiT [39] and SiT [40].
Models #Layers Hidden Size MLP #Heads Params
FAR-B 12 768 3072 12 130M
FAR-M 12 1024 4096 16 230M
FAR-L 24 1024 4096 16 457M
FAR-XL 28 1152 4608 18 674M
FAR-B-Long 12 768 3072 12 150M
FAR-M-Long 12 1024 4096 16 280M
3 Preliminary
3.1 Flow Matching
Flow Matching [41, 42, 43] is a simple alternative objective for training diffusion models. Rather than modeling the reverse process with stochastic differential equations, Flow Matching learns a continuous vector field that deterministically conntect two distribution.
Specifically, given a data sample $x_{0}\sim p_{\text{data}}(x)$ and a noise sample $x_{1}\sim\mathcal{N}(0,I)$, we construct a continuous trajectory connecting them via linear interpolation:
$x(t)=(1-t)x_{0}+tx_{1},\quad t\in[0,1].$ (1)
This formulation implies a constant velocity:
$\frac{dx(t)}{dt}=v^{*}=x_{1}-x_{0}.$ (2)
To enable the model to learn the optimal transport between the data and noise distributions, we introduce a learnable time-dependent velocity field $v_{\theta}(x,t)$. During training, a random time $t\sim U(0,1)$ is sampled, and the model is optimized by minimizing the following objective:
$\mathcal{L}(\theta)=\mathbb{E}_{x_{0},x_{1},t}\left[\left\|v_{\theta}(x(t),t)-v^{*}\right\|^{2}\right].$ (3)
Figure 3: Visualization of Attention Mask. FAR enables full attention within a frame while maintaining causality at the frame level. In long-context training, we adopt aggressive patchification for long-term context frames to reduce tokens.
3.2 Autoregressive Models
Autoregressive models are a class of probabilistic models where each element in a sequence is conditioned on its preceding elements, denote as context. Formally, given a sequence of tokens $(x_{1},x_{2},\dots,x_{n})$, an autoregressive model assumes that each token $x_{i}$ is generated based on its previous tokens $(x_{1},x_{2},\dots,x_{i-1})$. The generative process can be expressed as a factorization of the joint probability:
$p(x_{1},x_{2},\dots,x_{n})=\prod_{i=1}^{n}p(x_{i}|x_{1},x_{2},\dots,x_{i-1}).$ (4)
By modeling each token conditioned on its preceding tokens, autoregressive models naturally capture the sequential dependencies inherent in data.
Figure 4: Effect of Stochastic Clean Context. This technique eliminate training-inference gap in observed context.
4 FAR
In this section, we first present the framework of FAR in Sec. 4.1. Then, we discuss the difficulties and solutions in training FAR in Sec. 4.2. In Sec. 4.3, we analyze the key design that enables FAR for long-context video modeling.
4.1 Framework Overview
Architecture. As shown in Fig. 2 (a), FAR is built upon the diffusion transformer [40, 39]. We adopt the model configuration of DiT [39] and Latte [44], as listed in Table. I. The key architectural difference between FAR and video diffusion transformers (e.g., Latte [44]) lies in the attention mechanism. As shown in Fig. 3(a), for each frame, we apply causal attention at the frame level while maintaining full attention within each frame. We adopt this causal spatiotemporal attention for all layers, instead of the interleaved spatial and temporal attention used in Latte. In FAR, image generation and image-conditioned video generation are jointly learned thanks to the causal mask, whereas video diffusion transformer [44] requires additional image-video co-training.
Basic Training Pipeline. The training pipeline of FAR is illustrated in Fig. 2 (a). Given a video sequence $\mathbf{X}$, we first employ a pretrained VAE to compress it into the latent space $\mathbf{Z}\in\mathbb{R}^{T\times H\times W}$, where $T$, $H$, and $W$ denote the number of frames, height, and width of the latent features, respectively. Note that although we primarily adopt an image VAE in this work, FAR can also be trained with a video VAE since our autoregressive unit is the latent frame. Following diffusion forcing [17], we independently sample a timestep for each frame. We then interpolate between the clean latent and the sampled noise using Eq. Eq. 1 and apply the frame-wise flow matching objective in Eq. Eq. 3 for learning. The key difference between FAR and image flow matching lies in that we adopt causal spatiotemporal attention, allowing each frame to access previous context frames during denoising.
Figure 5: Comparison of FAR and video diffusion transformer. FAR achieves better convergence than video diffusion transformer in unconditional video generation on UCF-101.
Figure 6: KV Cache for Short-Video Modeling in FAR. We additionally add a caching step to encode current decoded frame into the KV cache for autoregressive generation.
Figure 7: Relation between Token Context Length and Vision Context Length. With the proposed long short-term context modeling, the token context length scales more slowly with increasing vision context length compared to uniform context modeling. When training on long videos, the reduced number of tokens leads to significantly lower training costs and memory usage.
4.2 Short-Video Modeling
Training-Inference Gap in Observed Context. As a hybrid AR-diffusion model, FAR also encounters a training-inference gap in the observed context. As illustrated in Fig. 2(a), each clean latent is fused with sampled noise for the flow matching objective, as defined in Eq. Eq. 1. Consequently, later frames can only access the noised version of previous frames during training. However, during inference, this leads to a distribution shift when clean context frames is used.
As shown in the example in Fig. 4(a), the training-inference gap in the observed context leads to a distribution shift when inferring with a clean context. Although adding mild noise to the context during inference can help mitigate this effect, it still causes low-level flickering, degrading the quality of the generated video. Recent works [20, 21] attempt to address this issue by maintaining a clean copy of the noised sequence during training. However, this approach doubles the training costs.
Our Solution: Stochastic Clean Context. To bridge the gap in observed context, we introduce stochastic clean context for training FAR. As illustrated in Fig. 2(a), we randomly replace a portion of the noised frames with their corresponding clean context and assign them a unique timestep embedding (e.g., -1) beyond the flow-matching timestep scheduler. These clean context frames are excluded from loss computation and are implicitly learned through later frames that use them as context. During inference, this unique timestep embedding guides the model to use clean context effectively.
Training FAR with stochastic clean context does not add extra computation and does not conflict with different timestep sampling strategies during training (e.g., logit-normal sampling [45]). It effectively resolves the training-inference discrepancy, as exemplified in Fig. 4(b).
FAR vs. Video Diffusion Transformer. FAR and video diffusion transformer differ only in their training schemes. FAR is trained with independent noise and causal attention, while the video diffusion transformer is trained with uniform noise and full attention. This raises an interesting question: Can FAR surpass video diffusion transformers? To explore this, we convert FAR to video diffusion transformer as a baseline, denoted as Video DiT. We align the training settings to compare the two paradigms. As shown in Fig. 5, FAR achieves better convergence than the Video DiT, demonstrating its potential to become a strong baseline for autoregressive video modeling.
Figure 8: Multi-Level KV Cache for Long-Context Video Modeling in FAR. When a frame leaves the short-term context window, we encode it to the L2 cache and re-encode the L1 cache in the window. We then use those encoded KV cache for decoding the current frame. Note that we divide the process into three steps for better illustration, though it can be merged into a single forward pass in implementation.
4.3 Long-Context Video Modeling
Token Redundancy in Long Video. Visual data contains spatial redundancy, causing vision tokens to expand much faster than language tokens as context increases. For example, a video sequence of 128 frames requires more than 8K tokens, as illustrated in Fig. 7(a). Consequently, training on long videos becomes computationally prohibitive, as shown in Fig. 7(b, c).
Our Solution: Long Short-Term Context Modeling. To address the token redundancy in video, we introduce long short-term context modeling, which exploits the spatial and temporal locality in video data. As illustrated in Fig. 2(b), we maintain a high-resolution short-term context window to learn fine-grained temporal consistency and a low-resolution long-term context window, where we adopt aggressive patchification to reduce the number of context tokens. During training, given that the data has a maximum sequence length of $m$ frames, we fix the short-term context window to $n$ frames and randomly sample the long-term context frames from the range $[0,m-n]$. The attention mask with long short-term context modeling is shown in Fig. 3(b), where the long-term context uses fewer tokens. As demonstrated in Fig. 7(a), this strategy ensures that increasing the vision context length maintains a manageable token context length. With long short-term context modeling, we can reduce the cost and memory usage of the long-video training significantly, as shown in Fig. 7(b, c). To prevent interference between long-term and short-term contexts, we adopt separate projection layers for each context, inspired by MM-DiT [45]. This approach results in a slightly larger number of parameters, referred to as FAR-Long in Table. I.
TABLE II: Experimental Configurations of FAR. We follow the evaluation settings from Latte [44], MCVD [46], and TECO [38].
Hyperparameters Short-Video Generation Short-Video Prediction Long-Video Prediction
Cond. UCF-101 Uncond. UCF-101 BAIR UCF-101 Minecraft DMLab
Dataset Configuration
Resolution 256/128 256/128 64 64 128 64
Total Training Samples 13,320 13,320 43,264 9,624 194,051 39,375
Training Configuration
Batch Size 32 32 32 32 32 32
Latent Size 8$\times$8 (DC-AE [47]) 8$\times$8 (DC-AE [47]) 8$\times$8 (DC-AE [47]) 8$\times$8 (DC-AE [47]) 8$\times$8 (DC-AE [47]) 8$\times$8 (DC-AE [47])
Training Sequence Length 16 16 32 16 300 300
LR $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
LR Schedule constant constant constant constant constant constant
Warmup Steps - - - - 10K 10K
Total Training Steps 400K 400K 200K 200K 1M 1M
Stochastic Clean Context 0.1 0.1 0.1 0.1 0.1 0.1
Short-Term Context Window 16 16 32 16 16 16
Long-Term Context Resolution - - - - 2$\times$2 2$\times$2
Evaluation Configuration
Samples 4$\times$2048 4$\times$2048 100$\times$256 100$\times$256 4$\times$256 4$\times$256
Guidance Scale 2.0 - - - 1.5 1.5
Reference Work Latte [44] Latte [44] MCVD [46] MCVD [46] TECO [38] TECO [38]
4.4 Inference-Time KV Cache
KV Cache for Short-Video Modeling. Due to the autoregressive nature of FAR, we can leverage KV-Cache to accelerate inference. As illustrated in Fig. 6, for each frame, we first use the flow-matching schedule to decode it into the clean latent frame. We then introduce an additional caching step to encode the clean latent frame into the KV cache. As discussed in Sec. 4.2, we use timestep $t=-1$ to denote the clean context frame in the caching step. These KV caches are subsequently used for autoregressive decoding of subsequent frames.
Multi-Level KV Cache for Long-Video Modeling. In long-context video modeling, we employ long short-term context to reduce redundant visual tokens. To accommodate this, we introduce a multi-level KV cache. As illustrated in Fig. 8, the frames in long-term context window is encoded into L2 cache (4 tokens per frame), while the frames in short-term context window is encoded into L1 cache. When decoding current frame but exceed the short-term context window, the earliest frame in the short-term context window is moved to the long-term context window and encode it into the L2 cache. Since this modifies the cache state, we subsequently re-encode the L1 cache of the frames in the short-term context window. The encoded cache is then used to decode the current frame. Note that in practice, these three steps can be merged into a single forward pass for efficiency.
5 Experiment
5.1 Implementation Details
We follow the DiT’s structure [39] to implement FAR. To compress video latents, we train a series of image DC-AE [47] on the corresponding dataset, resulting in 64 tokens per frame. All models are trained from scratch without image pretraining. We provide detailed training hyperparameters and evaluation setting in Table. II.
5.2 Quantitative Comparison
5.2.1 Video Generation
Dataset and Evaluation Settings. We benchmark both unconditional and conditional video generation on the UCF-101 dataset [48], which consists of approximately 13,000 videos. Following Latte [44], we use the entire dataset for training. For evaluation, we randomly sample 2,048 videos to compute FVD [49] against the ground-truth videos. For conditional video generation, we set the guidance scale to 2.0 during inference.
Main Results. From the results listed in Table. III, we achieve state-of-the-art performance in both unconditional and conditional video generation. Specifically, Latte [44] is based on video diffusion transformer, while OmniTokenizer [50] is based on Token AR. Our method significantly outperforms both. Furthermore, compared to recent frame-autoregressive models [20, 21], which require twice the training cost, FAR achieves superior performance without any additional training cost.
[…]
Methods Type Params Double Cond. Gen Uncond. Gen
Train Cost FVD2048 $\downarrow$ FVD2048 $\downarrow$
Resolution-128$\times$128
[…]
TATS [51] Token-AR 331 M 332 420
[…]
Resolution-256$\times$256
LVDM [14] Video-DiT 437 M - 372
Latte [44] Video-DiT 674 M - 478
CogVideo [27] Token-AR 9.4 B 626 -
OmniTokenizer [50] Token-AR 650 M 191 -
ACDIT [20] Frame-AR 677 M 111 -
MAGI [21] Frame-AR 850 M - 421
FAR-L (Ours) Frame-AR 457 M 113 303
FAR-XL (Ours) Frame-AR 674 M 108 279
TABLE IV: Quantitative Comparison on Short Video Prediction. We follow the evaluation setup of MCVD [52] and ExtDM [46], where $c$ denotes the number of context frames and $p$ denotes the number of predicted frames.
Methods Params $c=4,\ p=12$
[…]
RaMViD [53] 235 M 0.639 21.37 0.090 396.7
LFDM [54] 108 M 0.627 20.92 0.098 698.2
MCVD-cp [52] 565 M 0.658 21.82 0.088 468.1
ExtDM-K2 [46] 119 M 0.754 23.89 0.056 394.1
FAR-B (Ours) 130 M 0.818 25.64 0.037 194.1
(a) Evaluation on UCF-101 (64$\times$64)
Methods Params $c=2,\ p=14$ $c=2,\ p=28$
[…]
RaMViD [53] 235 M 0.758 17.55 0.085 166.5 0.691 16.51 0.109 238.7
LFDM [54] 108 M 0.770 17.45 0.084 167.6 0.730 16.68 0.106 276.8
VIDM [55] 194 M 0.763 16.97 0.080 131.7 0.728 16.20 0.096 194.6
MCVD-cp [52] 565 M 0.838 19.10 0.075 87.8 0.797 17.70 0.078 119.0
ExtDM-K4 [46] 121 M 0.845 20.04 0.053 81.6 0.814 18.74 0.069 102.8
FAR-B (Ours) 130 M 0.849 20.87 0.038 99.3 0.819 19.40 0.049 144.3
(b) Evaluation on BAIR (64$\times$64)
TABLE V: Quantitative Comparison on Long-Context Video Prediction. We follow the evaluation setup of TECO [38], where $c$ denotes the number of context frames and $p$ denotes the number of predicted frames.
Methods Params $c=144,\ p=156$ $c=36,\ p=264$
[…]
FitVid [56] 165 M 0.356 12.0 0.491 176
CW-VAE [57] 111 M 0.372 12.6 0.465 125
Perceiver AR [58] 30 M 0.304 11.2 0.487 96
Latent FDM [59] 31 M 0.588 17.8 0.222 181
TECO [38] 169 M 0.703 21.9 0.157 48
FAR-B-Long (Ours) 150 M 0.687 22.3 0.104 64
(a) Evaluation on DMLab (64$\times$64)
Methods Params $c=144,\ p=156$ $c=36,\ p=264$
[…]
FitVid [56] 176 M 0.343 13.0 0.519 956
CW-VAE [57] 140 M 0.338 13.4 0.441 397
Perceiver AR [58] 166 M 0.323 13.2 0.441 76
Latent FDM [59] 33 M 0.349 13.4 0.429 167
TECO [38] 274 M 0.381 15.4 0.340 116
FAR-M-Long (Ours) 280 M 0.448 16.9 0.251 39
(b) Evaluation on Minecraft (128$\times$128)
5.2.2 Short-Video Prediction
Dataset and Evaluation Settings. We evaluate FAR on the UCF-101 [48] and BAIR [60] datasets, following the evaluation settings in MCVD [52] and ExtDM [46]. We randomly sample 256 videos based on provided context frames, each with 100 different trajectories, and select the best trajectory to compute pixel-wise metrics. For FVD, we report the average over all trajectories.
Main Results. We summarize the results in Table. IV. Unlike previous works such as MCVD [52] and ExtDM [46], which introduce complex multi-scale fusion strategies and optical flow, FAR achieves superior results on both datasets without requiring additional design.
5.2.3 Long-Video Prediction
Dataset and Evaluation Settings. We benchmark long-context video modeling results on action-conditioned video prediction using the Minecraft and DMLab datasets [38]. The Minecraft dataset contains approximately 200K videos, while the DMLab dataset contains about 40K videos. Each video consists of 300 frames with action annotations. We follow the evaluation setup in TECO [38], which uses 144 observed context frames to predict 156 future frames and compute pixel metrics. Additionally, we compute FVD on 264 generated frames based on 36 context frames.
Main Results. We summarize the results in Table. V. The previous work, TECO [38], adopts aggressive downscaling for all frames to reduce tokens for temporal modeling, creating a trade-off between training efficiency and prediction accuracy. In contrast, FAR employs long short-term context modeling, effectively achieving the lowest prediction error (i.e., LPIPS) without prohibitive computation cost.
5.3 Qualitative Comparison
We present a qualitative comparison of long-video prediction in Fig. 10. Compared to previous methods, FAR effectively utilizes the observed context and generates predictions that most closely resemble the ground truth, demonstrating its ability to leverage long-range context.
Figure 9: Ablation Study of the KV Cache. FAR-Long with proposed multi-level KV cache achieves the best speedup on long videos.
Figure 10: Qualitative Comparison of Long-Context Video Prediction on DMLab. FAR fully utilizes the long-range context (144 frames), resulting in more consistent prediction (156 frames) compared to previous methods.
5.4 Ablation Study
Stochastic Clean Context. We have visualized the effectiveness of stochastic clean context in Fig. 4. Based on the quantitative evaluation of video prediction in Table. VI, FAR with stochastic clean context achieves significantly improved performance.
Multi-Level KV Cache. We evaluate the inference speed of FAR in Fig. 9. FAR without the KV cache incurs the highest inference cost, while the KV cache significantly reduces it. For long videos, FAR-Long with proposed multi-level KV cache achieves the lowest inference cost.
Long-Term Context Resolution. We investigate the impact of long-term context resolution on prediction accuracy and inference speed. As the context resolution increases, pixel-level metrics improve; however, the overall video quality remains similar. Nonetheless, inference speed significantly degrades at higher context resolutions due to the increased number of tokens involved in computation. Therefore, we select a 2$\times$2 resolution for the long-term context as a balance between computational efficiency and long-term context performance.
Short-Term Context Window Size. We evaluate the impact of the short-term context window size on performance. As shown in Fig. 11, video quality (FVD) quickly saturates as the short-term context window size increases, while pixel-level metrics continue to improve but also approach saturation at a window size of 16. Therefore, we set the short-term context window size to 16 by default.
TABLE VI: Ablation Study of Stochastic Clean Context on UCF-101. Stochastic clean context mitigates the training-inference discrepancy in observed context, leading to improved performance.
Methods $c=1,\ p=15$
[…]
FAR w/o. Stochastic Clean Context 0.540 16.42 0.211 399
FAR w/. Stochastic Clean Context 0.596 18.46 0.187 347
TABLE VII: Ablation Study on the Resolution of Long-Term Context. The speed is averaged over the generation of 300 frames.
[…]
$1\times 1$ 0.411 15.25 0.312 40 2.94
$2\times 2$ 0.423 15.84 0.291 40 2.88
$4\times 4$ 0.433 16.26 0.276 37 1.42
Figure 11: Ablation Study of the Short-Term Context Window Size. Performance saturates as the window size increases.
6 Conclusion
In this paper, we introduce FAR for long-context video modeling. FAR effectively learns causal dependencies across continuous frames and demonstrates faster convergence compared to video diffusion transformers. To address visual redundancy in long-video training, we further propose a long short-term context modeling strategy, which encodes both remote and nearby visual contexts using asymmetric patchification. This approach enables a manageable token length as context frames increase, reducing both time and memory costs during long-video training. FAR achieves state-of-the-art performance on both short- and long-video modeling tasks, underscoring its potential as a new foundation model for video generation.
Limitations. The primary limitation lies in the lack of scaled-up experiments. Although FAR demonstrates great potential, we still lack large-scale training on text-to-video generation datasets. Additionally, restricted by the available datasets, we only experiment with FAR on up to 300 frames (about 20 seconds), not fully investigating its ability on minute-level videos.
Future Work. One future direction is to scale up FAR and benchmark it against video diffusion transformers on large-scale text-to-video generation tasks. Additionally, we plan to simulate a longer video dataset (on the minute level) to better evaluate the model’s long-context capabilities. Finally, it would be interesting to explore whether FAR’s long-context modeling can enable video-level in-context learning.